# What can we learn from quantum convolutional neural networks?

### v2

@@ V Discussion
Another important question regarding the power of QCNNs and possible embeddings arises when we consider extensions from simple one-dimensional models and go into two dimensions. Namely, every feature can be in principle embedded into ground states of 2D spin models with distinct topological properties. In this case, the entanglement structure, correlations, and nature of phase transition change. And so does a critical exponent of the transition [90]. For instance, considering the toric code model or models with fractional statistics, one can enrich the basis for describing critical phenomena [104]. Can we match them with corresponding behaviour in hydrodynamical systems, thus ensuring quantum “simulation” of phase transitions in fluid dynamics? Can we extend it to chaotic systems? The development of QCNN understanding to embeddings with long-range topological order may well offer capabilities in classification and regression beyond currently explored states with limited entanglement.
@@ VI Conclusion
In this work we have developed a better understanding of quantum convolutional neural networks, specifically highlighting why they are so successful in quantum phase recognition. We have shown that supplied quantum states (features) can be understood in terms of hidden feature maps — quantum processes that prepare states depending on classical parameters $\bm{x}$. During the mapping process, we observed that the ground state preparation supplies a very beneficial basis set, which allows the building of a nonlinear quantum model for classification with decision boundaries that are sharp and generalize from only few data points and measurement samples. We show that single-qubit observables in QCNNs can be used to “pick up” the most beneficial basis functions, while leading to sampling advantage. The developed understanding also opens another perspective on quantum sensing aided by convolution-based quantum processing. Finally, motivated by classification, we applied the QCNN-based workflow with cluster model ground state embedding to solve problems in fluid dynamics (viscous Burgers equation) and wave physics (damped oscillator). This suggests quantum convolutional neural networks as a provably trainable tool for data-driven modelling of critical phenomena with quantum computers.
@@ Acknowledgements.
We thank Lachlan Lindoy and Ivan Rungger for useful discussion on the subject, and acknowledge the funding from UK EPSRC (award EP/Y005090/1).
@@ Appendix A Ground state preparation
using an appropriate adiabatic preparation path for Hamiltonian ^$\hat{\mathcal{H}}(x)$ (here $\mathcal{T}$ denotes time-ordering) [105]. This however depends on the properties of the Hamiltonian and gap closing.

### v3

@@ V Discussion
Another important question regarding the power of QCNNs and possible embeddings arises when we consider extensions from simple one-dimensional models and go into two dimensions. Namely, every feature can be in principle embedded into ground states of 2D spin models with distinct topological properties. In this case, the entanglement structure, correlations, and nature of phase transition change. And so does a critical exponent of the transition [90]. For instance, considering the toric code model or models with fractional statistics, one can enrich the basis for describing critical phenomena [104]. Can we match them with corresponding behaviour in hydrodynamical systems, thus ensuring quantum “simulation” of phase transitions in fluid dynamics? Can we extend it to chaotic systems? The development of exotic embeddings with long-range topological order may well offer capabilities in classification and regression beyond currently explored states with limited entanglement.
@@
We must also acknowledge the recent works by Cerezo et al. [105] and Bermejo et al. [106]. These works detail how some of trainable parameterized quantum circuits, in particular QCNNs, generally live in a polynomial subspace of operators. This opens up the possibility of their efficient classical treatment in the Heisenberg picture [77]. At the same time, working with non-trivial input states and specifically quantum data represents a beneficial scenario. Additionally we note that these studies suggest a possibility for QCNNs either being initialized or wandering out of this polynomial subspace during training. In this situation, full quantum simulation of the QCNN is necessary. Maintaining the trainability guarantees is also important, hence the entanglement of these quantum states must be limited. Finding such datasets is an open challenge.
@@ VI Conclusion
In this work we have developed a better understanding of working with quantum data. Dissecting a workflow of quantum convolutional neural networks as an example, we identify reasons for their success in tasks involving quantum data, particularly quantum phase recognition. We have shown that supplied quantum states (features) can be understood in terms of hidden feature maps — quantum processes that prepare states depending on classical parameters $\bm{x}$. During the mapping process, we observed that the ground state preparation supplies a very beneficial basis set, which allows the building of a nonlinear quantum model for classification with decision boundaries that are sharp and generalize from only few data points and measurement samples. We show that single-qubit observables in QCNNs can be used to “pick up” the most beneficial basis functions. This can lead to advantage in using a small number of samples, as compared to measure-first approaches with shadow tomography. The developed understanding also opens another perspective on quantum sensing aided by convolution-based quantum processing. Finally, motivated by classification, we applied the QCNN-based workflow with cluster model ground state embedding to solve problems in fluid dynamics (viscous Burgers equation) and wave physics (damped oscillator). This suggests quantum convolutional neural networks as a provably trainable tool for data-driven modeling of critical phenomena with quantum computers.
@@ Acknowledgements.
We thank Lachlan Lindoy and Ivan Rungger for useful discussion on the subject. O.K. and C.U. and acknowledge the funding from UK EPSRC (award EP/Y005090/1), as well as support from NATO SPS project MYP.G5860.
@@ Appendix A Ground state preparation
using an appropriate adiabatic preparation path for Hamiltonian ^$\hat{\mathcal{H}}(x)$ (here $\mathcal{T}$ denotes time-ordering) [107]. This however depends on the properties of the Hamiltonian and gap closing. The adiabatic approach also allows quantum models based on ground-state feature maps to be analysed as Fourier models [101].
@@
The ground state preparation strategy will determine the quantum resources (circuit depth, ancilla qubits) required to implement the ground state feature map, as well as the quality of the quantum model; the smooth basis functions visualised in Fig. 3(b) become noisy around the critical point [101].
@@ Appendix B Fixed QCNN analysis
Appendix C QCNN with arbitrary variational ansatz
In our previous studies we used a variational ansatz inspired by the fixed QCNN structure; this architecture proved to be extremely successful in recovering an accurate decision boundary for the cluster Hamiltonian. One question may arise; can we obtain similar success if the variational ansatz is chosen more arbitrarily, without guidance from the fixed (non-variational) QCNN structure. To investigate this, we select a more general variational ansatz, taken from [111]. Specifically, we select arbitrary SU(4) gates acting on pairs of neighbouring qubits to make up the convolutional layers, and controlled rotations as the pooling layers (still maintaining the 9-3-1 structure and translational invariance).
We train the model in the same way as before, and we compare the performance of the GS embedding QCNN with the guided and arbitrary variational ansatze in Fig. 8. We find that $100\%$ test accuracy is still achieved with this arbitrary ansatz trained on 4 training points, and the sharpness at criticality is maintained. Again, the importance of the specific basis of the GS embedding is on show.
Figure 8: Decision boundaries for the QCNN with GS embedding with the original, ”guided” ansatz detailed in Sec. IV.1, and an arbitrary variational ansatz.
Appendix D QCNN trained with shot noise
In Sec. IV.3, we demonstrated that sampling from the QCNN with GS embedding would generally lead to more accurate results, as the high variance of the Fourier embedding model induces errors in the finite shot regime. Here, we extend the analysis by both training and testing the model with a fixed number of shots. We apply the same procedure as before, just this time we train (and test) the model with 200 shots per expectation value evaluation.
Figure 9: Decision boundaries for the QCNN with GS and Fourier embeddings, trained and tested with 200 shots per expectation value estimate. Dots represent the test data points: the errors in the critical regime for the Fourier embedding are apparent.
As Fig. 9 shows, the QCNN model with GS embedding is still more effective in this regime. Although the tails of the GS feature map boundary are particularly noisy and do not quite match the exact boundary, the smooth, sharp transition at criticality is still maintained in the presence of shot noise. Meanwhile, the decision boundary for the QCNN with Fourier embedding is noisy throughout. Most importantly, the shot noise induces classification errors for data points around the critical region; only 97% test accuracy, compared to 100% for the GS embedding.